OpenStudy (anonymous):

Multiply (3)(−4)(2). −48 −24 24 48

OpenStudy (alexandervonhumboldt2):

-(3*4*2)

OpenStudy (anonymous):

@24

OpenStudy (alexandervonhumboldt2):

-24

OpenStudy (anonymous):

Multiply two and four-fifths negative five and two-thirds . negative ten and eight-fifteenths negative fifteen and thirteen-fifteenths ten and eight-fifteenths fifteen and thirteen-fifteenths

OpenStudy (anonymous):

please help

OpenStudy (alexandervonhumboldt2):

2*4/5=8/5 -5*2/3=-10/3

OpenStudy (alexandervonhumboldt2):

8/5*-10/3=-80/15

OpenStudy (anonymous):

B